This is the correct email address to response to:
Solve quad equation. 2x^2-4x=2x

Answer by algebrahouse.com(1659)

2x² - 4x = 2x
2x² - 6x = 0 {subtracted 2x from both sides}
2x(x - 3) = 0 {factored out the greatest common factor, 2x}
2x = 0 or x - 3 = 0 {set each factor equal to 0}
x = 0 or x = 3 {solved each equation for x}
www.algebrahouse.com